Find the particular solution for the following equation y"-y'-12y=te^4t


1
Expert's answer
2022-03-31T05:23:09-0400

"y''-y'-12y=te^{4t}\\\\\\lambda ^2-\\lambda -12=0\\Rightarrow \\lambda \\in \\left\\{ 4,-3 \\right\\} \\\\Thus\\,\\,the\\,\\,form\\,\\,of\\,\\,particular\\,\\,solution\\,\\,is\\,\\,tP\\left( t \\right) e^{4t},P\\,\\,is\\,\\,a\\,\\,polynomial\\,\\,of\\,\\,power\\,\\,1\\\\y=t\\left( A+Bt \\right) e^{4t}=\\left( At+Bt^2 \\right) e^{4t}\\\\y'=\\left( A+\\left( 2B+4A \\right) t+4Bt^2 \\right) e^{4t}\\\\y''=\\left( 8A+2B+\\left( 16A+16B \\right) t+16Bt^2 \\right) e^{4t}\\\\\\left( 8A+2B+\\left( 16A+16B \\right) t+16Bt^2-A-\\left( 2B+4A \\right) t-4Bt^2-12At-12Bt^2 \\right) e^{4t}=te^{4t}\\\\\\left\\{ \\begin{array}{c}\t7A+2B=0\\\\\t14B=1\\\\\t0B=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tA=-\\frac{1}{49}\\\\\tB=\\frac{1}{14}\\\\\\end{array} \\right. \\\\y=\\left( -\\frac{1}{49}t+\\frac{1}{14}t^2 \\right) e^{4t}\\\\"